It takes 36 sec for 0.01 moles of He to effuse through a pinhole in an effusion


apparatus. It takes 72 sec for 0.01 moles of another gas to effuse through the


same system. What is the molecular weight of the unknown gas?

1
Expert's answer
2022-02-14T13:56:37-0500

Solution;

From Graham's law of diffusion;

"\frac{t_{He}}{t_x}=\frac{n_{He}}{n_x}\sqrt{\frac{M_{He}}{M_x}}"

By direct substitution;

"\frac{36}{72}=\frac{0.01}{0.01}\sqrt{\frac{4}{M_x}}"

Simplify;

"(\frac12)^2=\frac{4}{M_x}"

"M_x=4\u00d74=16g\/mol"